OpenStudy (anonymous):

How do you solve the rational equation 1/4 + 1/8 = 1/x?

OpenStudy (anonymous):

does it = 8/3?

OpenStudy (anonymous):

@Nurali

OpenStudy (kaylaprincess):

no

OpenStudy (anonymous):

what does it = then?

OpenStudy (anonymous):

yes it does

OpenStudy (kaylaprincess):

lemme check 1 sec srry if im wrong

OpenStudy (kaylaprincess):

OOPS OMG SORRY :d

OpenStudy (kaylaprincess):

ha.. ha.

OpenStudy (helder_edwin):

\[\large \frac{1}{4}+\frac{1}{8}=\frac{1}{x} \] \[\large \frac{2+1}{8}=\frac{1}{x} \] \[\large \frac{3}{8}=\frac{1}{x} \] \[\large x=\frac{8}{3} \]